OpenStudy (anonymous):

solve the system of equations below: x-2y=3 2x-3y=9

OpenStudy (anonymous):

Hello would u like some help in very good at math I'm in 7th grade and I'm Cecilia

OpenStudy (anonymous):

@hartnn can u help me? :)

OpenStudy (anonymous):

i suck at these lol

hartnn (hartnn):

sure, lets try elimination method

hartnn (hartnn):

in the first equation you have x while in the 2nd one you have 2x

hartnn (hartnn):

so multiply entire 1st equation by 2 what do u get ?

hartnn (hartnn):

\(2 [x-2y] = 2 \times 3 \\ ... \)

OpenStudy (anonymous):

x=2y+3/2 ???

hartnn (hartnn):

did you try to isolate x ?

hartnn (hartnn):

that you need to do when you're using substitution method

hartnn (hartnn):

if you're more comfortable with substitution mentod, we can take that up

OpenStudy (anonymous):

i suck at math, so either way is going to be a challenge for me :(

hartnn (hartnn):

in the last question, you used substitution method, so in this one let us use elimination :)

hartnn (hartnn):

so multiply 2 throughout the equation x-2y =3 let me give you an example : Multiplying 3 throughout 4x+9y =-2 3[4x+9y] = -2*3 so we get 12x +27y = - 6 similarly what will u get if you multiply x-2y =3 by 2

OpenStudy (anonymous):

y=1/2x-3

OpenStudy (anonymous):

?

hartnn (hartnn):

ok, so you isolated y now :) lets go with substitution method only! yes, thats correct \(y = (1/2)(x-3)\) thats what you meant, right ?

OpenStudy (anonymous):

ooops yea :)

hartnn (hartnn):

now plug this 'y' in 2nd equation

hartnn (hartnn):

\(2x - 3 (1/2) (x-3) = 9 \) did you get that ? ^

hartnn (hartnn):

i just replaced y by 1/2 (x-3)

OpenStudy (anonymous):

so do i solve for x?

OpenStudy (anonymous):

x=9?

OpenStudy (anonymous):

(9,3)

hartnn (hartnn):

yes, you solve for x and x=9 is absolutely correct! :D

hartnn (hartnn):

yep, 9,3 :)

OpenStudy (anonymous):

:D thank you!!!!!

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

am i a super saiyan god now?? :) @hartnn lol

hartnn (hartnn):

lol, you'll become a god, if you could solve all math problems :P

OpenStudy (anonymous):

ah poo that'll never happen haha

OpenStudy (anonymous):

thats why i will forever just be bulma...

hartnn (hartnn):

don't lose hope, with practice its possible :)

hartnn (hartnn):

and you're a cuter version of bulma :)

OpenStudy (anonymous):

:3

OpenStudy (anonymous):

See your here too lol